# Bidirectional Consistency Models

### v2

@@ Abstract
Diffusion models (DMs) are capable of generating remarkably high-quality samples by iteratively denoising a random vector, a process that corresponds to moving along the probability flow ordinary differential equation (PF ODE). Interestingly, DMs can also invert an input image to noise by moving backward along the PF ODE, a key operation for downstream tasks such as interpolation and image editing. However, the iterative nature of this process restricts its speed, hindering its broader application. Recently, Consistency Models (CMs) have emerged to address this challenge by approximating the integral of the PF ODE, largely reducing the number of iterations. Yet, the absence of an explicit ODE solver complicates the inversion process. To resolve this, we introduce the Bidirectional Consistency Model (BCM), which learns a single neural network that enables both forward and backward traversal along the PF ODE, efficiently unifying generation and inversion tasks within one framework. Notably, our proposed method enables one-step generation and inversion while also allowing the use of additional steps to enhance generation quality or reduce reconstruction error. Furthermore, by leveraging our model’s bidirectional consistency, we introduce a sampling strategy that can enhance FID while preserving the generated image content. We further showcase our model’s capabilities in several downstream tasks, such as interpolation and inpainting, and present demonstrations of potential applications, including blind restoration of compressed images and defending black-box adversarial attacks.
@@ 1 Introduction
Two key components in image generation and manipulation are generation and its inversion. Specifically, generation aims to learn a mapping from simple noise distributions, such as Gaussian, to complex ones, like the distribution encompassing all real-world images. In contrast, inversion seeks to find the reverse mapping, transforming real data back into the corresponding noise  [^1] In some contexts, the inversion problem refers to restoring a high-quality image from its degraded version. However, in this paper, we define this term more narrowly as the task of finding the corresponding noise for an input image.. Recent breakthroughs in deep generative models [1, 2, 3, 4, 5, 6, 7] have significantly revolutionized this field. These models have not only achieved remarkable success in synthesizing high-fidelity samples across various modalities [8, 9, 10, 11], but have also proven effective in downstream applications, such as image editing, by leveraging the inversion tasks [12, 13, 14].
Particularly, score-based diffusion models (DMs) [4, 5, 6, 7, 15] have stood out among their counterparts for generation [16]. Starting from random initial noise, DMs progressively remove noise through a process akin to a numerical ODE solver operating over the probability flow ordinary differential equation (PF ODE), as outlined by Song et al. [7]. However, it typically takes hundreds of iterations to produce high-quality generation results. This slow generation process limits broader applications.
This issue has been recently addressed by consistency models (CMs) [17, 18] which directly computes the integral from any given time step to the PF ODE trajectory’s starting point. Building on top of CMs, Kim et al. [19] introduced the Consistency Trajectory Model (CTM), which estimates the integral between any two time steps along the trajectory towards the denoising direction. Through these approaches, the consistency model family enables image generation with only a single Number of Function Evaluation (NFE, i.e., number of network forward passes) while also offering a trade-off between speed and quality.
On the other hand, unfortunately, the inversion tasks remain challenging for DMs. First, the generation process in many DMs [5, 15] is stochastic and hence is non-invertible; second, even for methods that employ a deterministic sampling process [6], inverting the ODE solver relies on the assumption of local linearization, necessitating more iterations for a small reconstruction error; third, although CMs and CTM accelerate generation by learning the integral directly, this integration is strongly non-linear, making the inversion process even harder.
@@
    we train a single neural network that enables both forward and backward traversal (i.e., integration) along the PF ODE, unifying the generation and inversion tasks into one framework. Importantly, our training approach does not rely on pre-trained models or the use of LPIPS [20], ensuring the flexibility of the methods and the reliability of the results;
@@
    we demonstrate that BCMs can generate images or invert a given image with a single NFE, and can achieve improved sample quality or lower reconstruction error by chaining multiple time steps;
@@
    leveraging BCM’s reversibility, we introduce a sampling strategy that combines the benefits of introducing fresh noise in each sampling step while preserving high determinacy;
@@
    we apply BCM for image interpolation, inpainting, blind restoration of compressed image, and defending black-box adversarial attack, showcasing its potential in downstream tasks.
@@ 2 Background and Preliminary
Figure 1: An illustrative comparison of score-based diffusion models, consistency models, consistency trajectory models, and our proposed bidirectional consistency models. (a) DM estimates the score function at a given time step; (b) CM enforces self-consistency that different points on the same trajectory map to the same initial points; (c) CTM strengthens this principle of consistency. It can map any point at time $t$ back to another point at time $u\leq t$ in the same trajectory. (d) BCM is designed to map any two points on the same trajectory to each other, removing any restrictions on the mapping direction. When the mapping direction aligns with the diffusion direction, the model adds noise to an input image. Conversely, if the mapping direction is opposite, the model performs denoising. This approach unifies generation and inversion tasks into a single, cohesive framework.
@@
Score-based Diffusion Models. Score-based Diffusion Models (DMs) sample from the target data distribution by by progressively removing noise from a random ${\mathbf{x}}_{T}\sim N(0,T^{2}{\bm{I}})$. To achieve this, DMs first diffuse the data distribution $p_{\text{data}}({\mathbf{x}})$ through a stochastic differential equation (SDE) [7]:
$d{\mathbf{x}}_{t}=\bm{\mu}({\mathbf{x}}_{t},t)dt+\sigma(t)d{\mathbf{w}}_{t},$ (1)
which has a reverse-time SDE, as described by Anderson [21]:
$d{\mathbf{x}}_{t}=\left[\bm{\mu}({\mathbf{x}}_{t},t)-\frac{1}{2}\sigma^{2}(t)\nabla\log p_{t}({\mathbf{x}}_{t})\right]dt+\sigma(t)d\bar{{\mathbf{w}}}_{t},$ (2)
where $t\in[\varepsilon,T]$, $p_{t}$ is the marginal density of ${\mathbf{x}}_{t}$, and ${\mathbf{w}}_{t}$ and $\bar{{\mathbf{w}}}_{t}$ represents the standard Wiener process in forward and reverse time respectively. Note, that $p_{0}$ is our desired data distribution $p_{\text{data}}$. Remarkably, Song et al. [7] showed that there exists an ordinary differential equation, dubbed the Probability Flow (PF) ODE, whose solution trajectories have the same marginal density $p_{t}$ at time $t$:
$\frac{d{\mathbf{x}}_{t}}{dt}=\bm{\mu}({\mathbf{x}}_{t},t)-\frac{1}{2}\sigma^{2}(t)\nabla\log p_{t}({\mathbf{x}}_{t}).$ (3)
During training, DMs learn to estimate $\nabla\log p_{t}({\mathbf{x}}_{t})$ with a score model $\bm{s}({\mathbf{x}}_{t},t)$ with score matching [22, 7, 15]. And during sampling, DMs solve the empirical PF ODE from time $T$ to $\varepsilon$ with a numerical ODE solver. Following Karras et al. [15], we set $\mu=0,\sigma=\sqrt{2t},\varepsilon=0.002,T=80$, which yields the following empirical PF ODE:
$\hat{\frac{d{\mathbf{x}}_{t}}{dt}}=-t\bm{s}({\mathbf{x}}_{t},t).$ (4)
Consistency Models, Consistency Training, and improved Consistency Training. However, while the above procedure has proven effective in generating high-quality samples, the numerical ODE solver typically requires iterative evaluations of the network $\bm{s}({\mathbf{x}}_{t},t)$ and hence bottlenecks the generation speed. To this end, Song et al. [17] proposed Consistency Models (CMs) that train a network to estimate the solution to PF ODE directly, i.e.,
@@
The network $\bm{f}_{\bm{\theta}}(\cdot,\cdot)$ can be either trained by distillation from a pre-trained DM, or from scratch with Consistency Training (CT). Here, we describe the consistency training in more details since it lays the foundation of our proposed method: to begin, consistency training first discretizes the time horizon $[\varepsilon,T]$ into $N-1$ sub-intervals, with boundaries $\varepsilon=t_{1}<t_{2}<\dots<t_{N}=T$. And the training objective is defined as
@@
$\bm{\theta}$ and $\bm{\theta}^{-}$ represents the parameters of the online network and a target network, respectively. The target network is obtained by $\bm{\theta}^{-}\leftarrow\texttt{stopgrad}\left(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta}\right)$ at each iteration. $\lambda(\cdot)$ is a reweighting function, ${\mathbf{x}}$ represents the training data sample, and ${\mathbf{z}}\sim\mathcal{N}(0,{\bm{I}})$ is a random Gaussian noise. During training, $N$ is gradually increased, allowing the model to learn self-consistency in an incremental manner. Additionally, Song et al. [17] proposed to keep track of an exponential moving average (EMA) of the online parameter $\bm{\theta}$. Specifically, after optimizing $\bm{\theta}$ in each iteration, CMs update the EMA by
$\bm{\theta}_{\text{EMA}}\leftarrow\mu_{\text{EMA}}\bm{\theta}_{\text{EMA}}+(1-\mu_{\text{EMA}})\bm{\theta}.$ (7)
After training, CMs discard the online parameters and generate samples with $\bm{\theta}_{\text{EMA}}$.
Besides, in a follow-up work, Song and Dhariwal [18] suggested to use the Pseudo-Huber loss for $d$, along with other techniques that include setting $\mu=0$ (i.e., $\bm{\theta}^{-}\leftarrow\texttt{stopgrad}\left(\bm{\theta}\right)$), proposing a better scheduler function for $N$, adapting a better reweighting function $\lambda(t_{n})=\nicefrac{{1}}{{|t_{n}-t_{n+1}|}}$. Dubbed improved Consistency Training (iCT), these modifications significantly improve the performance. Therefore, unless otherwise stated, we inherit these improving techniques in our work.
Consistency Trajectory Model. While CMs learn the integral from an arbitrary starting time to 00, Consistency Trajectory Model (CTM) [19] learns the integral between any two time steps along the PF ODE trajectory towards the denoising direction. More specifically, CTM learns
@@
CTM demonstrates that it is possible to learn a stronger consistency than that induced by CMs: two points ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{u}$ along the same trajectory not only can map back to the same initial point ${\mathbf{x}}_{0}$, but also can map from ${\mathbf{x}}_{t}$ to ${\mathbf{x}}_{u}$, provided $u\leq t$. This inspires us for a stronger consistency with a bijection between ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{u}$.
However, it is worth noting that, while sharing some motivation, our methodology is foundationally different from that proposed in CTM. CTM is based on a completely different network parameterization, whereas ours is a natural extension of the approach used by EDM [15]. Moreover, CTM largely relies on adversarial loss [1], while our proposed method, following Song and Dhariwal [18], is free from LPIPS [20] and adversarial training.
@@ 3 Methods
To this end, we adjust the network parameterization and the training objective of consistency training [17, 18]. These modifications are detailed in Sections 3.1 and 3.2. Besides, we introduce new sampling schemes capitalizing on our model’s invertibility, which are elaborated upon in Section 3.3. Finally, we end this section by presenting the inversion algorithm of BCM in Section 3.4.
Algorithm 1 Bidirectional Consistency Training (differences between BCT and CT/iCT [17, 18] are highlighted in Red $\blacksquare$)
@@
      Sample $n^{\prime}\sim\tilde{p}(n^{\prime}|N(k))$, where $\tilde{p}(n^{\prime}|N(k))\propto\begin{cases}0,&\text{if }n^{\prime}=n,\\ p(n^{\prime}|N(k)),&\text{otherwise.}\end{cases}$
@@
      $\mathcal{L}(\bm{\theta})\leftarrow\mathcal{L}_{\text{CT}}(\bm{\theta})+\mathcal{L}_{\text{ST}}(\bm{\theta})$;
     $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$;
@@
     $\bm{\theta}_{\text{EMA}}\leftarrow\mu_{\text{EMA}}\bm{\theta}_{\text{EMA}}+(1-\mu_{\text{EMA}})\bm{\theta}$;
     $k\leftarrow k+1$;
@@ 3.1 Network Parameterization
We first describe our network parameterization. Our network takes in three arguments: 1) the sample ${\mathbf{x}}_{t}$ at time $t$ along the PF ODE, 2) the time step $t$, and 3) the target time step $u$, and outputs the sample at time $u$, i.e., ${\mathbf{x}}_{u}$. To achieve this, we directly expand the models used in Consistency Models (CMs) [17, 18] with an extra argument $u$. In CMs, the networks first calculate Fourier embeddings [23] or positional embeddings [24] for the time step $t$, followed by two dense layers. Here, we simply concatenate the embeddings of $t$ and $u$, and double the dimensionality of the dense layers correspondingly.
Also, following CMs [17, 18] and EDM [15], instead of directly learning $\bm{f}_{\bm{\theta}}$, we train $\bm{F}_{\bm{\theta}}$ and let:
Based on an extension of EDM [15], we set
@@
Notice that $c_{\text{skip}}(t,t)=1$ and $c_{\text{out}}(t,t)=0$, which explicitly enforce the boundary condition $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t},t,t)={\mathbf{x}}_{t}$. We include detailed derivations in Appendix B.

### v4

@@ Abstract
Diffusion models (DMs) are capable of generating remarkably high-quality samples by iteratively denoising a random vector, a process that corresponds to moving along the probability flow ordinary differential equation (PF ODE). Interestingly, DMs can also invert an input image to noise by moving backward along the PF ODE, a key operation for downstream tasks such as interpolation and image editing. However, the iterative nature of this process restricts its speed, hindering its broader application. Recently, Consistency Models (CMs) have emerged to address this challenge by approximating the integral of the PF ODE, largely reducing the number of iterations. Yet, the absence of an explicit ODE solver complicates the inversion process. To resolve this, we introduce Bidirectional Consistency Model (BCM), which learns a single neural network that enables both forward and backward traversal along the PF ODE, efficiently unifying generation and inversion tasks within one framework. We can train BCM from scratch or tune it using a pretrained consistency model, which reduces the training cost and increases scalability. We demonstrate that BCM enables one-step generation and inversion while also allowing the use of additional steps to enhance generation quality or reduce reconstruction error. We further showcase BCM’s capability in downstream tasks, such as interpolation and inpainting. Our code and weights are available at https://github.com/Mosasaur5526/BCM-iCT-torch.
@@ 1 Introduction
Two key components in image generation and manipulation are generation and its inversion. Specifically, generation aims to learn a mapping from simple noise distributions, such as Gaussian, to complex ones, like the distribution encompassing all real-world images. In contrast, inversion seeks to find the reverse mapping, transforming real data back into the corresponding noise  [^1] The inversion problem also refers to restoring a high-quality image from its degraded version. However, in this paper, we define it more narrowly as the task of finding the corresponding noise for an input image.. Recent breakthroughs in deep generative models (Goodfellow et al., 2014; Kingma & Welling, 2013; Dinh et al., 2014; Song & Ermon, 2019; Ho et al., 2020; Song et al., 2021a; b) have revolutionized this field. These models have not only achieved remarkable success in synthesizing high-fidelity samples across various modalities (Karras et al., 2021; Rombach et al., 2022; Kong et al., 2021; OpenAI, 2024), but have proven effective in downstream applications, such as image editing, by leveraging the inversion process (Mokady et al., 2023; Huberman-Spiegelglas et al., 2023; Hertz et al., 2023).
Particularly, score-based diffusion models (DMs) (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2021a; b; Karras et al., 2022) have stood out for generation (Dhariwal & Nichol, 2021). Starting from random initial noise, DMs progressively remove noise through a process akin to a numerical ODE solver operating over the probability flow ordinary differential equation (PF ODE), as outlined by Song et al. (2021b). However, it typically takes hundreds of iterations to produce high-quality generation results. This slow generation process limits broader applications.
This issue has been recently addressed by consistency models (CMs) (Song et al., 2023; Song & Dhariwal, 2024), which directly compute the integral of PF ODE trajectory from any time step to zero. Similar to CMs, Kim et al. (2024) introduced the Consistency Trajectory Model (CTM), which estimates the integral between any two time steps along the trajectory towards the denoising direction. Through these approaches, the consistency model family enables image generation with only a single Number of Function Evaluation (NFE, i.e., number of network forward passes) while offering a trade-off between speed and quality.
On the other hand, unfortunately, the inversion tasks remain challenging for DMs. First, the generation process in many DMs (Ho et al., 2020; Karras et al., 2022) is stochastic and hence is non-invertible; second, even for methods that employ a deterministic sampling process (Song et al., 2021a), inverting the ODE solver necessitates at least hundreds of iterations for a small reconstruction error; third, although CMs and CTM accelerate generation by learning the integral directly, this integration is strongly non-linear, making the inversion process even harder.
@@
    We train a single neural network that enables both forward and backward traversal (i.e., integration) along the PF ODE, unifying the generation and inversion tasks into one framework. We can either train the model from scratch or fine-tune it from a pretrained consistency model to reduce training cost and increase scalability.
@@
    We demonstrate BCM can generate images or invert a given image with a single NFE, and can achieve improved sample quality or lower reconstruction error by chaining multiple time steps.
@@
    Leveraging BCM’s capability to navigate both forward and backward along the PF ODE trajectory, we introduce two sampling schemes and a combined approach that has empirically demonstrated superior performance.
@@
    We apply BCM for image interpolation, inpainting, and blind restoration of compressed images, showcasing its potential versatile downstream tasks.
@@ 2 Background and Preliminary
Figure 1: An illustrative comparison of score-based diffusion models, consistency models, consistency trajectory models, and our proposed bidirectional consistency models. (a) DM estimates the score function at a given time step; (b) CM enforces self-consistency that different points on the same trajectory map to the same initial points; (c) CTM strengthens this principle of consistency, which maps a point at time $t$ back to another point at time $u\leq t$ along the same trajectory. (d) BCM is designed to map any two points on the same trajectory to each other, removing any restrictions on the mapping direction. When the mapping direction aligns with the diffusion direction, the model adds noise to an input image. Conversely, if the mapping direction is opposite, the model performs denoising. This approach unifies generation and inversion tasks into a single, cohesive framework.
@@
Score-based Diffusion Models. Score-based Diffusion Models (DMs, Song et al., 2021b) sample from the target data distribution by progressively removing noise from a random ${\mathbf{x}}_{T}\sim N(0,T^{2}{\bm{I}})$. To achieve this, DMs first diffuse the data distribution $p_{\text{data}}({\mathbf{x}})$ through a stochastic differential equation (SDE): $d{\mathbf{x}}_{t}=\bm{\mu}({\mathbf{x}}_{t},t)dt+\sigma(t)d{\mathbf{w}}_{t},$ which has a reverse-time SDE, as described by Anderson (1982): $d{\mathbf{x}}_{t}=\left[\bm{\mu}({\mathbf{x}}_{t},t)-\sigma^{2}(t)\nabla\log p_{t}({\mathbf{x}}_{t})\right]dt+\sigma(t)d\bar{{\mathbf{w}}}_{t},$ where $t\in[0,T]$ [^2] To avoid numerical issues, we always set $t$ in $[0.002,T]$ in practice. However, to keep the notation simple, we will ignore this small initial value $0.002$ when describing our methods in this paper. , $p_{t}$ is the marginal density of ${\mathbf{x}}_{t}$, and ${\mathbf{w}}_{t}$ and $\bar{{\mathbf{w}}}_{t}$ represents the standard Wiener process in forward and reverse time respectively. Note, that $p_{0}$ is our desired data distribution $p_{\text{data}}$. Remarkably, Song et al. (2021b) showed that there exists an ordinary differential equation, dubbed the Probability Flow (PF) ODE, whose solution trajectories have the same marginal density $p_{t}$ at time $t$:
${d{\mathbf{x}}_{t}}/{dt}=\bm{\mu}({\mathbf{x}}_{t},t)-1/2\sigma^{2}(t)\nabla\log p_{t}({\mathbf{x}}_{t}).$ (1)
During training, DMs learn to estimate $\nabla\log p_{t}({\mathbf{x}}_{t})$ with a score model $\bm{s}({\mathbf{x}}_{t},t)$ with score matching (Hyvärinen & Dayan, 2005; Song et al., 2021b; Karras et al., 2022). And during sampling, DMs solve the empirical PF ODE from time $T$ to 00 with a numerical ODE solver. Following Karras et al. (2022), we set $\mu=0,\sigma=\sqrt{2t},T=80$.
Consistency Models, Consistency Training, and improved Consistency Training. However, while the above procedure has proven effective in generating high-quality samples, the numerical ODE solver typically requires iterative evaluations of the network $\bm{s}({\mathbf{x}}_{t},t)$ and hence bottlenecks the generation speed. To this end, Song et al. (2023) proposed Consistency Models (CMs) that train a network to estimate the solution to PF ODE directly, i.e.,
@@
The network $\bm{f}_{\bm{\theta}}(\cdot,\cdot)$ can be either trained by distillation or from scratch with Consistency Training (CT). Here, we describe the consistency training in more detail since it lays the foundation of our proposed method: to begin, consistency training first discretizes the time horizon $[0,T]$ into $N-1$ sub-intervals, with boundaries $0=t_{1}<t_{2}<\dots<t_{N}=T$. The training objective is defined as
@@
$\bm{\theta}$ and $\bm{\theta}^{-}$ represents the parameters of the online network and a target network, respectively. The target network is obtained by $\bm{\theta}^{-}\leftarrow\texttt{stopgrad}\left(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta}\right)$ at each iteration. $\lambda(\cdot)$ is a reweighting function, ${\mathbf{x}}$ represents the training data sample, and ${\mathbf{z}}\sim\mathcal{N}(0,{\bm{I}})$ is a random Gaussian noise. During training, $N$ is gradually increased, allowing the model to learn self-consistency in an incremental manner. Additionally, Song et al. (2023) proposed to keep track of an exponential moving average (EMA) of the online parameter $\bm{\theta}$. Specifically, after optimizing $\bm{\theta}$ in each iteration, CMs update the EMA by $\bm{\theta}_{\text{EMA}}\leftarrow\mu_{\text{EMA}}\bm{\theta}_{\text{EMA}}+(1-\mu_{\text{EMA}})\bm{\theta}.$ After training, CMs discard the online parameters and generate samples with $\bm{\theta}_{\text{EMA}}$.
Besides, in a follow-up work, Song & Dhariwal (2024) suggested to use the Pseudo-Huber loss for $d$, along with other techniques that include setting $\mu=0$ (i.e., $\bm{\theta}^{-}\leftarrow\texttt{stopgrad}\left(\bm{\theta}\right)$), proposing a better scheduler function for $N$, adapting a better reweighting function $\lambda(t_{n})=\nicefrac{{1}}{{|t_{n}-t_{n+1}|}}$. Dubbed improved Consistency Training (iCT), these modifications significantly improve the performance. Therefore, unless otherwise stated, we inherit these improving techniques in our work.
Consistency Trajectory Model. While CMs learn the integral from an arbitrary starting time to 00, Consistency Trajectory Model (CTM) (Kim et al., 2024) learns the integral between any two time steps along the PF ODE trajectory towards the denoising direction. More specifically, CTM learns
@@
CTM demonstrates that it is possible to learn a stronger consistency: two points ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{u}$ along the same trajectory not only can map back to the same initial point ${\mathbf{x}}_{0}$, but also can map from ${\mathbf{x}}_{t}$ to ${\mathbf{x}}_{u}$, provided $u\leq t$. This inspires us for a stronger consistency with a bijection between ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{u}$.
However, it is worth noting that, while sharing some motivation, our methodology is foundationally different from that in CTM. Our network adopts a different parameterization, which is a natural extension of that by EDM (Karras et al., 2022). Moreover, CTM largely relies on adversarial loss (Goodfellow et al., 2014), while our proposed method, following Song & Dhariwal (2024), is free from LPIPS (Zhang et al., 2018) and adversarial training.
@@ 3 Methods
Algorithm 1 Bidirectional Consistency Training (Red indicates differences from CT/iCT (Song et al., 2023; Song & Dhariwal, 2024)
@@
      Sample $n^{\prime}\sim\tilde{p}(n^{\prime}|N(k))$, where $\tilde{p}(n^{\prime}|N(k))\propto\text{$\begin{cases}0,&\text{if }n^{\prime}=n,\\ p(n^{\prime}|N(k)),&\text{otherwise.}\end{cases}$ }$
@@
      $\mathcal{L}_{\text{BCT}}(\bm{\theta})\leftarrow\mathcal{L}_{\text{CT}}(\bm{\theta})+\mathcal{L}_{\text{ST}}(\bm{\theta})$;
     $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}_{\text{BCT}}(\bm{\theta})$;
@@
     $\bm{\theta}_{\text{EMA}}\leftarrow\mu_{\text{EMA}}\bm{\theta}_{\text{EMA}}+(1-\mu_{\text{EMA}})\bm{\theta},k\leftarrow k+1$;
@@ 3.1 Network Parameterization
We first describe the network parameterization. Our network takes in three arguments: 1) the sample ${\mathbf{x}}_{t}$ at time $t$, 2) the current time step $t$, and 3) the target time step $u$, and outputs the sample at time $u$, i.e., ${\mathbf{x}}_{u}$. To achieve this, we directly expand the models used in Consistency Models (CMs) (Song et al., 2023; Song & Dhariwal, 2024) with an extra argument $u$. In CMs, the networks first calculate Fourier embeddings (Tancik et al., 2020) or positional embeddings (Vaswani et al., 2017) for the time step $t$, followed by two dense layers. Here, we simply concatenate the embeddings of $t$ and $u$, and double the dimensionality of the dense layers correspondingly.
Similar to CMs (Song et al., 2023) and EDM (Karras et al., 2022), instead of directly learning $\bm{f}_{\bm{\theta}}$, we train $\bm{F}_{\bm{\theta}}$ and let $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t},t,u)=c_{\text{skip}}(t,u){\mathbf{x}}_{t}+c_{\text{out}}(t,u)F_{\bm{\theta}}(c_{\text{in}}(t,u){\mathbf{x}}_{t},t,u)$, where
@@
Note that $c_{\text{skip}}(t,t)=1$ and $c_{\text{out}}(t,t)=0$, which explicitly enforce the boundary condition $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t},t,t)={\mathbf{x}}_{t}$. We detail our motivation and derivations in Appendix D.